OpenStudy (anonymous):

[6.01] Select the system of equations that corresponds to the given graph.

OpenStudy (anonymous):

-2x - y = 4 2x + y = -2 -8x + 4y = -16 -4x + 2y = -8 x + 2y = 4 2x - 4y = 8 2x + 3y = 6 3x + 2y = 4

OpenStudy (anonymous):

???

jhonyy9 (jhonyy9):

from this graph you get that when x=4 y=0 and when x=0 y=-2 and when x=0 y=2

jhonyy9 (jhonyy9):

now you need check these roots to what equations coresponde,are right ok ?

OpenStudy (anonymous):

-2x - y = 4 2x + y = -2 -8x + 4y = -16 -4x + 2y = -8 x + 2y = 4 2x - 4y = 8 2x + 3y = 6 3x + 2y = 4 i can choose from these

OpenStudy (anonymous):

-2x - y = 4 2x + y = -2 -8x + 4y = -16 -4x + 2y = -8 x + 2y = 4 2x - 4y = 8 or 2x + 3y = 6 3x + 2y = 4

jhonyy9 (jhonyy9):

yes from these you need to chose it

OpenStudy (anonymous):

is it the 1st one?

jhonyy9 (jhonyy9):

are you sure ?

jhonyy9 (jhonyy9):

when y=0 rsulte x=4

jhonyy9 (jhonyy9):

1. will be x+2y=4

jhonyy9 (jhonyy9):

because here you see if you check it that for y=0 result x=4 yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

you need getting again two ok ?

jhonyy9 (jhonyy9):

for x=0 y=2 and for x=0 y=-2

OpenStudy (anonymous):

ok...?

jhonyy9 (jhonyy9):

the second one will be for y= -2

jhonyy9 (jhonyy9):

and for y=2 can you get it ?

jhonyy9 (jhonyy9):

x+2y=4

jhonyy9 (jhonyy9):

so and with this you have got all ok ?

jhonyy9 (jhonyy9):

do you understand it why are these equations right for these points ?

OpenStudy (anonymous):

yes but i dont get which choice is right

jhonyy9 (jhonyy9):

so from this graph you see there that for these intersection points of thses lines with axe x and y there are these points of intersection for x you see 4 and for y 2 and -2 yes ?

jhonyy9 (jhonyy9):

so after this you just need to get these equations for what has like roots these points of intersection 2,-2 and 4

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

so than good luck bye

OpenStudy (dan815):

you can just solve both of the line equations y= mx + b

OpenStudy (dan815):

use the points (0,2) and (4,0) and (4,0) and (0,-2) so for the first y=mx+b you get y = (rise/run) x + b y= ((0-2)/(4-0))x+b y=-1/2x+b now substitute the point ( 4,0) and get the b (0)=-1/2(4)+b) b = 2 so first equation is y = -2x + 2 or y+1/2x=2 or 2y+x=4 you can similarly solve for the 2nd equation and get y= 1/2(x) - 2 or 2y-x=-4 or -4y+2x=8 if u multiply by -2

OpenStudy (dan815):

when i mean or thats just multiplying both sides of the equation by a constant so theyre all equal

OpenStudy (dan815):

i hope that makes sense